OpenStudy (anonymous):

lim (x-1)/ [(x+3)^1/2 -2] x->1

OpenStudy (anonymous):

\[{x-1} \over {\sqrt{x+3}-2}\]

OpenStudy (anonymous):

rationalize the denominator

OpenStudy (anonymous):

\[\lim (x-1)\times (\sqrt{x+3} +2)/x+3-4= \sqrt{x+3}+2=\]4